a car accelarates from 0/ms to 30/ms in 5 second has an accelaration of...........m/s2?
1
Expert's answer
2012-10-22T11:26:15-0400
Let a be the acceleration.
Then the velocity v changes in time t by the
formula
v = at
hence
a = v/t = 30/5 = 6 m/s^2